OpenStudy (anonymous):

How to convert 6e^(-ipi/3) to polar form?

OpenStudy (anonymous):

I haven't a clue how to start, so I'd just like a push in the right direction if you could.

OpenStudy (anonymous):

Opps, make that standard form.

jimthompson5910 (jim_thompson5910):

6e^(-i*pi/3) is in the form r*e^(i*theta) where r = 6 theta = pi/3

jimthompson5910 (jim_thompson5910):

so the polar form is simply (r, theta) = (6, pi/3) you can express this in trig form to get r*cis(theta) = 6*cis(pi/3) recall that cis(theta) = cos(theta) + i*sin(theta)

OpenStudy (freckles):

\[ae^{i \theta}=a(e^{i \theta})=a(\cos(\theta)+i \sin(\theta)) =acos(\theta)+i a \sin(\theta)\]

OpenStudy (freckles):

I actually thought it was already in polar form

OpenStudy (freckles):

doing this other way would convert to rectangular form

jimthompson5910 (jim_thompson5910):

http://tutorial.math.lamar.edu/Extras/ComplexPrimer/Forms.aspx based on that, polar form is z = r*[cos(theta) + i*sin(theta)]

jimthompson5910 (jim_thompson5910):

he refers to r*e^(i*theta) as "Exponential Form"

OpenStudy (freckles):

and also according to that site e^(i theta) is the exponential polar form so it is still polar form

OpenStudy (anonymous):

Thanks, my teacher didn't discuss this at all.

OpenStudy (freckles):

or maybe according to that site they are calling that form just exponential form lol